OpenStudy (anonymous):

Find the fifth term and the nth term of the geometric sequence whose initial term, a, and common ratio, r, are given. a = √ 5 ; r = √ 5 answers: a5=5√ 5 ; 5^n/2 a5=25√ 5 ; 5^n/2 a5=5√ 5 ; 5^n-1 a5=25√ 5 ; 5^n

OpenStudy (campbell_st):

a term in a geometric series is found usng \[T _{n}=a \times r^{n-1}\] in your question you have a and r and n = 5 substitute and evaluate

OpenStudy (anonymous):

\[a_1=\sqrt{5}, a_2=\sqrt{5}\times \sqrt{5}=5,a_3=5\sqrt{5}, a_4=25,a_5=25\sqrt{5}...\]

OpenStudy (anonymous):

the answer es 25√5 but wich one is the nth term?????

OpenStudy (anonymous):

nth term is \(\sqrt{5}\times \sqrt{5}^{n-1}=\sqrt{5}^n=5^{\frac{n}{2}}\)

OpenStudy (anonymous):

5^(5/2)

OpenStudy (anonymous):

yes, that one

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw ( you did it before i did)

OpenStudy (campbell_st):

then go back to the formula and replace use n... where you used 5